OpenStudy (anonymous):

What is 14 percent of 35?

OpenStudy (anonymous):

4.9

OpenStudy (anonymous):

4.9

OpenStudy (anonymous):

It's 35*.14=?

OpenStudy (anonymous):

NOT ALLOWED TO GIVE FULL ANSWERS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (itsonlycdeee):

14/100 * 35

OpenStudy (mathstudent55):

14 percent of 35 Translate the above sentence into math. "percent" means "divided by 100" "of" means "multiply" 14 percent of 35 = 14/100 * 35 = 0.14 * 35